OpenStudy (lena772):

Generate the first 5 terms of this sequence: f(1) = 0 and f(2) = 1, f(n) = f(n - 1) + f(n - 2), for n > 2. 0, -1, 1, 0, 2 0, 1, 1, 2, 3 0, 1, 2, 2, 3 0, 1, 1, 2, 2

OpenStudy (lena772):

@SACAPUNTAS @kraken123

OpenStudy (lena772):

@ranga

OpenStudy (lena772):

I don't know how to do this one.

OpenStudy (anonymous):

Sorry, I have to leave now. You are getting this, though... keep at it. :)

OpenStudy (lena772):

Thanks for all your help, g'bye.

OpenStudy (anonymous):

its Fibonacci series

OpenStudy (ranga):

f(n) = f(n - 1) + f(n - 2) Put n =3 f(3) = f(2) + f(1) = ?

OpenStudy (lena772):

I don't understand? :/

OpenStudy (anonymous):

every number after the 2 nd term is the sum of previous two numbers

OpenStudy (anonymous):

so the answer is b)

OpenStudy (ranga):

They tell you f(1) = 0 and f(2) = 1 and so what is f(3)? f(3) = f(2) + f(1) = ? It looks like there are others willing to give you the answers without helping you learn. So I am out.

OpenStudy (anonymous):

Here is an outline of what it would look like in python: http://www.codeskulptor.org/#user20_mKYol66NQV_1.py (Need Firefox or Chrome for best compatibility). Examine the code and see if you can work through it.

OpenStudy (lena772):

I actually want to learn though. @ranga Thanks for explaining.